OpenStudy (anonymous):

Simplify: 2-3i / 4-3i

OpenStudy (anonymous):

I know I need to factor but I can't find any common factors

OpenStudy (anonymous):

first you need to multiply the top and the bottom by 4+3i to get rid of the i's in the denominator. \[\frac{ 2-3i }{ 4 - 3i }* \frac{ 4+3i }{ 4+3i }\]

OpenStudy (anonymous):

i can do this even though there is no equality?

OpenStudy (anonymous):

just letting you know that i^2=-1

OpenStudy (anonymous):

yes, i am aware of this

OpenStudy (anonymous):

thanks for letting me know though. appreciate it.

OpenStudy (anonymous):

\[\frac{ 8 + 6i - 12i - 9i^2 }{ 16 + 12i - 12i - 9I^2 }\]

OpenStudy (anonymous):

\[\frac{ 17 - 6i}{ 25 }\]

OpenStudy (anonymous):

\[\frac{ 17 }{ 25 } - \frac{ 6 }{ 25 }i\]

OpenStudy (anonymous):

awesome. thanks for the help. I really should know all this, my brain has just been melted by too much borderlands. thanks man

OpenStudy (anonymous):

You're welcome. :)